The government of Broncoland uses monetary policy tools similar to the Federal Reserve System of the United States and defines its monetary aggregates the same way as the Federal Reserve System of the United States. The required reserve ratio in Broncoland is 10%. The following information also applies to the government of Broncoland:

Bank deposits at the central bank = $200 million

Currency held by the public = $150 million

Currency in bank vaults = $100 million

Checkable bank deposits = $500 million

Traveler’s checks = $10 million

With the information above in mind, answer the following questions:

What is the value of M1?

What is the value of required reserves?

Are commercial banks holding excess reserves? If yes, what is the value of excess reserves?

Can central banks increase checkable deposits? If yes, please explain how, and state by how much checkable deposits can increase.
Answer #1

M1 = Coins and currency in circulation + Checkable deposits + Travelers checks

M1 = 150 + 500+10 = 660 million

Required reserves = required reserve ratio x currency in bank vaults

Required reserves = 10% x [ Currency held in deposits and vaults + currency in Central bank]

Required reserves = 0.1x[100+500+200] = 80 million

Yes banks holding excess reserves

Excess reserves = Reserves - Required reserves

Excess reserves = 200 - 80 = 120 million